OpenStudy (anonymous):

SOMEONE PLEASE HELP with the restriction x ?2, what is the simplifed form for 2x-7/ x-2 minus x+5/x-2?

OpenStudy (anonymous):

Since the denominators are the same you can just subtract the tops (2x-7)-(x+5)=x-2/x-2=1